can someone show how to solve the system by elimination method. Show all steps and check the answer? Thanks!
x - 2y = -1
3X + y = 4

Found 2 solutions by stanbon, doukungfoo:
Answer by stanbon(48568)

1st: x - 2y = -1
2nd: 3x + y = 4
----------------------
Multiply 2nd thru by 2 to get:
3rd: 6x + 2y = 8
----------------
Add 1st and 3rd to get:
7x =7
x = 1
-------------
Substitute that into 1st to solve for Y:
1 -2y = -1
2y = 2
y = 1
------------------
Check the answer in 2nd:
3*1 + 1= 4
3 + 1 = 4
4 = 4
===============
Cheers,
Stan H.

Answer by doukungfoo(195)

In systems of equations where the coefficients of terms containing the same variable are opposites, the elimination method can be applied by adding the equations. If the coefficients of those terms are the same, the elimination method can be applied by subtracting the equations.
Note: the coeffiecient is the number being multiplied by the variable.

However in this system of equations there are no terms that are opposite.
But we can make the y terms opposite by multiplying the second equation by two
to get 6x+2y=8
now you can use the elimination method
x - 2y = -1
6x + 2y = 8

add x terms to get 7x
add y terms to get 0
add -1 +8 to get 7
now you have
7x=7
x=1
now just use that in either equation to solve for y